OpenStudy (anonymous):

Carlee purchased a motorcycle for $9,941. She made a down payment of $3,795. She applied for a four-year installment loan with an interest rate of 13.2%. What is the total cost of the motorcycle after four years?

OpenStudy (anonymous):

this interest rate is every year Right ?

OpenStudy (anonymous):

well as i understood the total cash coast =9941$ she Paid: 3795 at 1st & the rest is 4-year installment loan the rest=6146$ on 4 years so she will pay:1536.5 every year the interest rate=13.2% but i don't knw this rate % is from the whole coat=9941$ , or from the yearly interest=1536.5$ so i'll walk in the 2 ways & u decide which is right so1st way: since the rate is 13.2% from the whole coast=9941 so the Rate=(13.2X9941)/100=1312.212$ so we will time this rate By 4-years then add it to the rest coast of the cycle to get=(1312.212 X 4)+6146=11394.848$

OpenStudy (anonymous):

The 2nd way: Since the rate=13.2% from the interest loan of 1-year=1536.5$ so the rate=(13.2X1536.5)/100=207.4275$ so we'll time this rate By 4 & then add it to the rest of the cycle Coast to get=(207.4275X4)+6146=6975.71$

OpenStudy (anonymous):

LooK this the coast of the cycle after 4-years Neglecting the 1st Payment which=3795$

OpenStudy (anonymous):

i hope it worked ?!

OpenStudy (rockinhood):

What was the answer for these?